The current and future legal

framework

A Threat to Free Software and Computer

Science

Alexandre Dulaunoy

adulau@foo.be - adulau@freearchive.org

NGO/ASBL AEL Association Electronique Libre (Belgium)

Agenda 1/2

• General Principles around Freedom in Computer Science

• Freedom : an Ethical approach

• Legal Framework

• A Paradoxical Situation or the importance of the balance

• Legal Protection against the Circumvention of the Copy Protection

• The land of DMCA, EUCD, DMCRA...

Agenda 2/2

• Imposing the use of copyright protection systems • “Cybercrime” law

• Software patents

• Free Software developers/company

• Conclusion ?

• Bibliography/ref/link

Freedom in Computer Science 1/4

• Q&A • François-René Rideau

"The Laws of Human Beings must not contradict that which Nature renders impossible to contradict"

• The exclusion principle of physical objects

• The non-exclusion principle of non-physical objects

You can kill a human being but not an idea...

• Bernard Lang

• Ecology of ideas

"...des idées, des concepts, et des objets immatériels qui, pour se développer, s’enrichir et s’épanouir, ont besoin de circuler, de s’associer, de se confronter, de se fertiliser les uns les autres."

• John Perry Barlow (EFF) - 1996 (CDA reaction)

• A Declaration of the Independence of Cyberspace

"Your legal concepts of property, expression, identity, movement, and context do not apply to us. They are all based on matter, and there is no matter here."

• Richard Stallman (FSF)

• Four Freedoms of Free Software

1. The freedom to run the program, for any purpose.

2. The freedom to study how the program works, and adapt it to your needs.

3. The freedom to redistribute copies so you can help your neighbor.

4. The freedom to improve the program, and release your improvements to the public, so that the whole community benefits.

Accessibility of source code is a necessary condition for free software.

The current and future legal framework – p.7/20

Freedom : an important note

• Freedom is an ethical issue

• Freedom is not a legal issue

• Freedom is here without legislation

• Disclaimer :

• But after we’ll talk about Freedom and legislation because legislation is hurting Freedom...

Legal framework 1/2

• The balance of a legal framework that protects Freedom and a legal framework that removes Freedom. Or The

paradox of the legal blob.

• Legal framework that can be used to "protect" Freedom :

• The use of the Copyright (for example The GNU General Public License) (Berne Convention)

• The European Directive allows reverse engineering for interoperability ( Japan)

• Europe : Munich Convention (article 52c) (paradoxical usage)

Legal framework 2/2

• WIPO Copyright Treaty (1996 version) (author->editor) • Article 11 (–> local law : DMCA / EUCD)

• Article 12 (–> “DRM”)

• Software Patents (in US / Japan)

• International “Cybercrime” Treaty (Council of EU + ...)

• DRM law (proposal directive)

• Data retention (european directive proposal 351/133)

Legal Protection against the Circumvention of the Copy Protection

• WIPO Copyright Treaty (1996 version)

• Obligations concerning Technological Measures Contracting Parties shall provide adequate legal protection and effective legal remedies against the circumvention of effective technological me asures that are used by authors in connection with the exercise of their rights under this Treaty or the Berne Convention and that restrict acts, in respect of their works, which are not authorized by the authors concerned or

permitted by law.-> 2001/29/CE -> DMCA,EUCD

Legal Protection against the Circumvention of the Copy Protection

• (US) DMCA (Digital Millenium Copyright Act) • Multiple cases : bnetd, DeCSS, Elcomsoft, ...

• (EUROPE) EUCD (European Copyright Directive) (2001/29/CE)

• Application in the national law for November 2002 • Check http://www.eucd.info/

• -> Fair-use, Computer Science Research, Competition, Cryptography, ...

Imposing the use of copyright protection systems

• SSSCA Security System Standards and Certification Act proposal

• CBDTPA Consumer Broadband and Digital Television

Promotion Act

• A work in-progress exists on a proposal Directive for Europe on imposing DRM (-> slowdown)

• This type of law would require equipment manufacturers to embed government-approved copy protection systems into all computer equipment.

• On the technical side, there is also some evolution on

“DRM”...

The current and future legal framework – p.13/20

Issues on imposing the use of copyright protection systems

• Access to education • The Right to Read (Richard Stallman, 1996)

• Universal Declaration of Human Rights (article 27)

• Reverse engineering • Entrance price for RE will be higher (more difficult?)

• Life of information

• Open Format and long-time preservation of information

(archive)

• Privacy

Cybercrime law

• International Cybercrime Treaty has been signed on 23 november 2001. (Council of Europe, US, Canada, Japan ...)

1. It includes a list of crimes that each member country must have on its books

2. It requires each participating nation to grant new powers of search and seizure to its law enforcement authorities

3. It requires law enforcement in every participating country to assist police from other participating countries by cooperating

• A proposal European directive (COM2002 173final)

• Reverse engineering issue, Computer Security...

The current and future legal framework – p.15/20

Software patents

• Software patents are forbidden in Europe (following the Munich Convention)

• Software patents are legal in the US (and in Japan) (LPF - 1991)

• European Patent Office (EPO) is trying to pass software patents with some tricks

• A proposal Directive in favor of Sofware Patents in EU • Implementing idea -> the idea itself

• Viscosity of research (danger for the innovation)

• Free Software & FS industry (RF licensing / RAND)

The current and future legal framework – p.16/20

Activism (from People to Free Software companies)

• Create/use/modify/distribute Free Software

• Promote privacy & anonymity

• Promote Open Standards and Free of access Standards

• Boycott and talk about products including “DRM”

• Use correct terms (from GNU/Linux to not use "Intellectual Property"...)

• Join / Help / make a donation to an association (FSF, EFF, FSF Europe,...)

• ...

The current and future legal framework – p.17/20

Conclusion?

• Freedom is an ethical issue

• Long terms plan are difficult or impossible

• Day to day activism to promote the ethical value of Freedom

• “Freedom of the Information” exists but some are trying to create new boundaries...

Bibliography

Right to read, R. Stallman, http://www.gnu.org/philosophy/right-to-read.html Manifeste de la Libre information, http://fare.tunes.org/Manifeste.fr.html A Declaration of the Independence of Cyberspace http://www.eff.org/ barlow/Declaration-Final.html Cybercrime law http://www.aclu.org/issues/privacy/Cybercrime_Feature.html http://europa.eu.int/eur-lex/en/com/pdf/2002/com2002_0173en01.pdf Ten Fallacies of Internet Censorship (Prof. Jeffrey Shallit) http://www.csi.uottawa.ca/ dduchier/misc/infohighway1.html Software Patents http://www.ael.be/node.php?id=52 http://www.ffii.org/index.en.html

Q&A

• adulau@foo.be - alexandre.dulaunoy@ael.be

• 3B12 DCC2 82FA 2931 2F5B 709A 09E2 CD49 44E6 CBCD